ROTATIONAL ENERGY TRANSFER IN o-/p-H_{2}+HD AT LOW TEMPERATURES

\underline{RENAT A. SULTANOV}, DENNIS GUSTER, Business Computing Research Laboratory, St. Cloud State University, 31 Centennial Hall, 720 Fourth Avenue South, St. Cloud, MN 56301-4498.

Quantum-mechanical close-coupling calculations for rotational state resolved cross sections and thermal rate coefficients for the o-/p-H_{2}+HD collisions of astrophysical interest will be presented. Recently developed new global H_{2}-H_{2} potential energy surface^{a} has been appropriately adopted for H_{2} + HD and applied. The low temperature limit of o-/p-H_{2}+HD is investigated, which is of significant astrophysical interest in regard to the cooling of primordial gas and the interstellar media. A test of convergence and the results for cross sections with the new potential^{a} are obtained for a wide range of kinetic velocities including values down to ∼10 m/s. Sharp resonances have been reproduced in the cross sections of some transition states at very low energies. Our results revealed^{b}, that for lower quantum transition states the new surface provides cross sections very close to those obtained in previous works, where the authors adopted some old potentials for H_{2}-H_{2}. However, for higher quantum states we found significant disagreements with previous results. Additionally, in our calculations new resonances are calculated in the 1300±100 m/s region^{b}. The value of the resonances are relatively large, and it may exert a strong influence on the cooling processes in primodial gas and interstellar media.

aA.I. Boothroyd, P. G. Martin, W. J. Keogh, M. J. Peterson, J. Chem. Phys., 116, 666 (2002). bR.A. Sultanov, D. Guster, Chem. Phys. Lett. 436, 19 (2007).